From mar@roe.ac.uk Mon Jul 30 16:24:08 2012 Date: Mon, 30 Jul 2012 16:00:21 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Andrew Lawrence , Bob Mann , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Norman Gray , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives team meeting, 27 July 2012 Present: MAR, RSC, NJC, ETWS, MSH Apologies: JPE, AL, KTN, STV, RGM, NG, NCH, RPB DoNM: 10am Friday August 10th 2012 in the Vista Hut (subject to non-demolition) NCH in the chair. Actions discharged this time: ----------------------------- ACTION: NJC to clarify with RGMcM if he needs anything from WFAU for the current ESO deliverables NJC and MAR had been in contact with RGMcM. It seems that RGMcM is after additional header entries in the web UI FITS output e.g. TUCDi column cards, we will try and accommodate. Actions partly discharged but continuing: ----------------------------------------- ACTION: RSC to scope out the areas/situations where we'd like to test MonetDB i.e. where do the current bottlenecks lie? Continues; ACTION: ETWS (with help from NJC as necessary) to add FeII 1.644 micron filter details into the WSA. Continues; ETWS had made some code changes but further set-up required. Actions carried forward from 13/07/12 meeting --------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. ACTION: NCH to nudge JRL and ask if VDFS-like bulk slurps of GES data can be set up. Continues; (might have been done but NCH not present) ACTION: NJC to communicate with Nottingham concerning incorporation of UDS in DR10. Continues; ACTION: NJC to contact VVV team to see if they're happy with world release of DR1. Continues; Specific points and new actions: -------------------------------- Project management: The grant renewal case (Q2 2013 - Q1 2016) has been submitted. WFCAM, VISTA and VST updates: Nothing to report this time. Comments and issues arising from CASU minutes: No new minutes this time. Networking: Network transfers are working well. ETWS noted: -- All data transferred for VISTA (up to April), WFCAM (up to May) and OmegaCam (up to June). WSA/VSA/OSA/GES Operations: WSA - CU1 & 2'ed up to the end of May. VSA - Cu3'ed up to the end of April. ETWS noted: -- Changed non-survey names for VISTA to not use the observation number. OSA - All available data transferred. ETWS noted: -- Updated code for the OSA metadata ingests and monitoring. We will get illumination corrected catalogues from CASU and ingest these but will ingest uncorrected catalogues when they're not available. This will be flagged in the detection table as well as in ppErrBits. Hardware and Systems: MSH noted: The new NAS box was up and running. The migration of Khufu had been completed. Survey Data Release: WSA - awaiting QC (eyeball results and scripted) VIKING - awaiting QC for P87/P88 and CU4 ingests of remaining P88 data. VIDEO - ESO release currently running. VHS - next release to be initiated following CU4 ingests of remaining P88 data VMC - ESO release made and handed over to Maria-Rosa. QC done for next proprietary release which will be made following ingest of remaining P88 data. Software: RSC noted: "ESO's request for a catalogue of unique sources led to the [re-]discovery that a lot of entries in our [seamed] merged source table aren't necessarily unique for a variety of reasons. One particular case for VIDEO created a lot of head scratching that ultimately revealed a feature of the pairing algorithm's handshake process, whereby multiple [non-unique] sources could potentially be defined within the pairing criterion. This was inevitable and perhaps entirely intentional to ensure the best possible matching within a strict criterion, thus "non-unique" sources in the source table are both to be expected and a good thing (it merely illustrates the unavoidable uncertainty). NCH and RSC have been enhancing the GES database with the latest atomic line data from the GES Wiki whilst awaiting access to bulk transfer the latest processing of the full data." NJC reported working on finalise flat file ingest and ESO release code. Non-survey Data Release: None. Astrogrid deployment & Data Analysis services: MSH reported that the Astrogrid registry, CVS and Wiki services would be moving to ROE next week. Miscellaneous: NJC noted that the deadline for registration for the ESO's October workshop on "Science from the Next Generation Imaging and Spectroscopic Surveys" had been and gone. ACTION: NJC to contact ESO and see if we could still send someone to the ESO workshop in October. -- Scanned by iCritical.